25 bags of wheat each weighing 40 kg cost Rs. 2750. Find the cost of 35 bags of wheat if each
bag weighs 50 kg.

Grade:8

3 Answers

Saswat Pani
6 years ago
25 bags of wheat each weighing 40 kg cost Rs. 2750.Then 1 bag weighing 40 kgs costs 2750/25 =110rs1 bag containg 50 kg weight is 110/40*50 =142.50So 35 bags costs 4987.50rs
C.Vishaal
6 years ago
25 bags = 2750Rs.
1 bag = 2750/25 = 110Rs
 
So 1 bag of 40 kg costs 110 Rs.
 
1 bag of 50 kg = 110/40 x 50 = 142.50 Rs
35 bags of 50 kg will therefore cost 142.50 x 35 which gives you an answer of 4987.50Rs.
 
Hope this answer helps.
Regards
C.Vishaal
jegatha
5 years ago
25 bags=rs 2750
therefore 1 bag= 2750/25=rs.110
1 bag of 40 kg costs=rs110
then 1 bag of 50 kg costs=(110/40) x 50=rs.137.5
therefore cost of 35 bags of 50 kg=rs.137.5 x 35=rs.4812.5